OpenStudy (anonymous):

what is the next number in the sequence? 729-1331-2197-3375-4913

ganeshie8 (ganeshie8):

hint : 729 = 9^3 1331 = 11^3

OpenStudy (anonymous):

6859?

OpenStudy (anonymous):

19^3

ganeshie8 (ganeshie8):

Yes!

ganeshie8 (ganeshie8):

\[\large a_n = (2n+7)^3 \]

OpenStudy (anonymous):

now your getting cheeky! ;)

ganeshie8 (ganeshie8):

hmm